OpenStudy (ell):

Help Pwease!! Write the first three terms of the binomial expansion, expressing the result in simplified form. (x^2+2)^14

OpenStudy (p0sitr0n):

expansion: (n) a^i b^n-i (i)

OpenStudy (p0sitr0n):

(n) (i) is n! / (i! n-i !)

OpenStudy (p0sitr0n):

i! is i*i-1*i-2*i-3... for instance, 3! = 3*2*1=6